OpenStudy (anonymous):

Sasha buys two ribbons of equal length. She cuts 9 pieces of the same length, L, from the first ribbon and finds that 5 inches are left over. She then cuts two pieces of the same length, L, from the second ribbon and finds that 40 inches are left over. What was the length of each ribbon when she bought them?

OpenStudy (anonymous):

9L+5=2L+40 solve for L and find the ribbon length like 9L+5 or 2L+40

hero (hero):

R= the length of both ribbons L = the length of the pieces R - 9L = 5 R - 2L = 40 R = 5 + 9L R = 40 + 2L 5 + 9L = 40 + 2L 9L - 2L = 40 - 5 7L = 35 L = 35/7 L = 5 R = 5 + 9(5) R = 5 + 45 R = 50 Thus the length of the original ribbon is 50 inches.